TABLE 1 Future Value of $1 FV = $1(1+i) 20.0% 1.20000 1.44000 ndi 1.0% 1 1.01000 2 1.02010 3 1.03030 4 1.04060 5 1.05101 1.5TABLE 2 Present Value of $1 $1 PV (1+1) wi 1.0% 1 0.99010 2 0.98030 3 0.97059 4 0.96098 5 0.95147 1.5% 0.98522 0.97066 0.956TABLE 3 Future Value of an Ordinary Annuity of $1 EVA_ (1 + i) - 1 ni 1.0% 1.5% 2.0% 2.5% 3.0% 3.5% 4.0% 4.5% 5.0% 5.5% 6.0%TABLE 4 Present Value of an ordinary Annuity of $1 1- (+1) PVA== ni 1 2 3 4 5 1.0% 1.5% 0.99010 0.98522 1.97040 1.95588 2.940TABLE 5 Future Value of an Annuity Due of $1 (1+i) -11 FVAD - = *(1+i) wi 1.0% 1.5% 2.0% 2.5% 3.0% 3.5% 4.0% 4.5% 5.0% 5,5%TABLE 6 Present Value of an Annuity Due of $1 PVAD= |x (1+i) ni 1 2 3 4 5 1.0% 1.00000 1.99010 2.97040 3.94099 4.90197 1.5% 1On January 1, 2021, Ithaca Corp. purchases Cortland Inc. bonds that have a face value of $320,000. The Cortland bonds have a
Answer #1

Part 1

Bond fair value

$300061

Payments

Amount

Present value factor

=

Present value

Interest

14400

X

12.46221

=

179456

Principal

320000

X

0.37689

=

120605

Bond price

$300061

320000*9%/2 = 14400

PV of $1 @ i= 5%, n = 20 is 12.46221

PV ordinary annuity of $1 @ i= 5%, n = 20 is 0.37689

Part 2

Date

General journal

Debit

Credit

January 01, 2021

Investment in bonds

320000

Discount on bond investment

19939

Cash

300061

June 30, 2021

Cash

14400

Discount on bond investment

603

Interest revenue ((320000-19939)*10%/2)

15003

June 30, 2021

No journal entry required

December 31, 2021

Cash

14400

Discount on bond investment

573

Interest revenue ((320000-19939-603)*10%/2)

14973

Part 3

Date

General journal

Debit

Credit

January 01, 2021

Investment in bonds

320000

Discount on bond investment

19939

Cash

300061

June 30, 2021

Cash

14400

Discount on bond investment

603

Interest revenue ((320000-19939)*10%/2)

15003

June 30, 2021

Net unrealized holding gains and losses—I/S

18905

Fair value adjustment (300061+603-281759)

18905

December 31, 2021

Cash

14400

Discount on bond investment

573

Interest revenue ((320000-19939-603)*10%/2)

14973

December 31, 2021

Net unrealized holding gains and losses—I/S

17389

Fair value adjustment (300061+603+573-264943-18905)

17389

June 30

Payments

Amount

Present value factor

=

Present value

Interest

14400

X

11.95038

=

172085

Principal

320000

X

0.34273

=

109674

Bond price

$281759

320000*9%/2 = 14400

PV of $1 @ i= 5.5%, n = 20 is 11.95038

PV ordinary annuity of $1 @ i= 5.5%, n = 20 is 0.34273

December 31

Payments

Amount

Present value factor

=

Present value

Interest

14400

X

11.46992

=

165167

Principal

320000

X

0.31180

=

99776

Bond price

$264943

320000*9%/2 = 14400

PV of $1 @ i= 6%, n = 20 is 11.46992

PV ordinary annuity of $1 @ i= 6%, n = 20 is 0.31180